OpenStudy (anonymous):

7(2m-2)-5m=9m-14 is this identity contratiction or conditional?

OpenStudy (anonymous):

How far did you get?

OpenStudy (anonymous):

I have a hard time solving them I can identify them but I cant solve them

OpenStudy (anonymous):

So for a first step, it might be good to make it simpler. You'll want to distribute the first term 7(2m-2)

OpenStudy (anonymous):

5m?

OpenStudy (anonymous):

7m I mean?

OpenStudy (anonymous):

Whenever you have multiplication outside of parentheses, each term (in this case, 2m and -2) needs to be multiplied by whatever is on the outside (7 in this case)

OpenStudy (anonymous):

First, 7*2m = ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you do the m with the 2

OpenStudy (anonymous):

Right, to get 14. And then you have an m attached, so that gives you 7*2m = 14m

OpenStudy (anonymous):

Oops, sorry, misread

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

ok so.....????

OpenStudy (anonymous):

Then we do the same thing with 7 and -2 to get -14

OpenStudy (anonymous):

ok Im getting this.

OpenStudy (anonymous):

So we know that 7(2m - 2) is the same as 14m - 14. We're going to choose 14m - 14 because it's a little simpler. Now we have 14m - 14 - 5m = 9m - 14

OpenStudy (anonymous):

Okay so far?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Okay. So, if you see something like that in a future problem, the first step is probably to distribute. What's our next step?

OpenStudy (anonymous):

cominbe like terms?

OpenStudy (anonymous):

polpak,,,, hey u!

OpenStudy (anonymous):

Actually, just rewrite the equation as it is now and I think you'll be able to identify it.

OpenStudy (anonymous):

Also, send me an email so I have yours ;)

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

Which terms should you combine?

OpenStudy (anonymous):

I really don't think you need to do the combining part. If you write the whole equation you can identify it now.

OpenStudy (anonymous):

my comp froze sorry guys

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

I got it.

OpenStudy (anonymous):

Im going to run and grab my charger for my comp should I combine terms or no?

OpenStudy (anonymous):

Can you write the expression for just the left side of the equation now (after distributing)?

OpenStudy (anonymous):

then compare that with the right side.

OpenStudy (anonymous):

There are several ways to do it. I was suggesting one way.

OpenStudy (anonymous):

Oh, certainly.

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

Me or Polpak?

OpenStudy (anonymous):

doesnt matter you were helping me just fine.

OpenStudy (anonymous):

Okidoki. Which like-terms should you combine (and what do you get)?

OpenStudy (anonymous):

14m-5m=9m?

OpenStudy (anonymous):

Good :). Now what is your equation?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

What do you get for the equation after you combine the like-terms?

OpenStudy (anonymous):

identity?

OpenStudy (anonymous):

Yes :), perfect.

OpenStudy (anonymous):

but wait whaat about the -14 '''

OpenStudy (anonymous):

yay!!!!!!!!

OpenStudy (anonymous):

You have 9m-14=9m-14

OpenStudy (anonymous):

omg I cant believe I solved it thank you so much

OpenStudy (anonymous):

you made it pretty easy

OpenStudy (anonymous):

Thanks :). You did great. Algebra is difficult, but it just takes practice to get better.

OpenStudy (anonymous):

Look at these,

OpenStudy (anonymous):

The sum of a number and -14 is -32 . find the number

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

-14x is 32?

OpenStudy (anonymous):

What is a sum?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

Yes, so you'll have to have addition in there. I like the choice of a variable for the number you don't know.

OpenStudy (anonymous):

(-14x)+is(-32)

OpenStudy (anonymous):

How do we write "is" in math-speak?

OpenStudy (anonymous):

Would we solve the same way we did b4?

OpenStudy (anonymous):

*?

OpenStudy (anonymous):

Similarly, but first we have to get the equation.

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

We have a sum, so we don't have multiplication.

OpenStudy (anonymous):

Instead of multiplication, we should have addition.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-14x+ (-32)

OpenStudy (anonymous):

should ther be a equal sign after the -32?

OpenStudy (anonymous):

Closer. We have "The sum of a number and -14." This means that we are adding an unknown number (which you've been calling x) to -14. As far as the equal sign goes, we almost always translate "is" as "=" in word problems.

OpenStudy (anonymous):

Everything before the word "is": The sum of a number and -14 This means all that goes before the = sign. Everything after the word "is" will go on the right of the = sign.

OpenStudy (anonymous):

You're getting closer. Give it another try.